# Functional normalizing flow for statistical inverse problems of partial differential equations

### v1

@@ Abstract
Inverse problems of partial differential equations are ubiquitous across various scientific disciplines and can be formulated as statistical inference problems using Bayes’ theorem. To address large-scale problems, it is crucial to develop discretization-invariant algorithms, which can be achieved by formulating methods directly in infinite-dimensional space. We propose a novel normalizing flow based infinite-dimensional variational inference method (NF-iVI) to extract posterior information efficiently. Specifically, by introducing well-defined transformations, the prior in Bayes’ formula is transformed into post-transformed measures that approximate the true posterior. To circumvent the issue of mutually singular probability measures, we formulate general conditions for the employed transformations. As guiding principles, these conditions yield four concrete transformations. Additionally, to minimize computational demands, we have developed a conditional normalizing flow variant, termed CNF-iVI, which is adept at processing measurement data of varying dimensions while requiring minimal computational resources. We apply the proposed algorithms to two typical inverse problems governed by a simple smooth equation and the steady-state Darcy flow equation. Numerical results confirm our theoretical findings, illustrate the efficiency of our algorithms, and verify the discretization-invariant property.
@@ 1 Introduction
Driven by their widespread applications in seismic exploration, radar imaging, and other fields, inverse problems involving partial differential equations (PDEs) have witnessed significant advancements in recent decades [28, 2]. As computational power continues to grow, researchers are increasingly focused on not only obtaining estimated solutions but also conducting statistical analyses to quantify uncertainties, which is crucial for tasks such as artifact detection [50]. The Bayesian inverse approach offers a robust framework for addressing inverse problems within the context of PDEs by transforming them into statistical inference problems, thereby enabling the analysis of parameter uncertainties [17].
Typically, inverse problems of PDEs are posed in infinite-dimensional spaces [28], which presents challenges for applying well-established finite-dimensional Bayesian inference techniques [5, 27]. To bridge this gap, two predominant strategies have been developed:
@@
    Discretize-then-Bayesianize: PDEs are initially discretized to approximate the original problem in some finite-dimensional space, and the reduced, approximated problem is then solved using Bayes’ method [27].
@@
    Bayesianize-then-discretize: Bayes’ formula and algorithms are initially constructed in infinite-dimensional space, and after the infinite-dimensional algorithm is built, some finite-dimensional approximation is carried out [17].
Both approaches offer distinct advantages and disadvantages. The Discretize-then-Bayesianize approach allows us to employ all the Bayesian inference methods developed in the statistical literature [27, 12] to solve the inverse problems. However, due to the infinite-dimensional nature of the original problems, two critical challenges arise:
@@
    Model consistency: Finite-dimensional models, such as those representing prior probability measures, exhibit distinct properties compared to their infinite-dimensional counterparts. A notable example is the total variation prior, which has been extensively studied in [32, 16].
@@
    Algorithm applicability: To preserve the intrinsic structure of original infinite-dimensional problems, algorithms initially designed for finite-dimensional spaces must be carefully adapted to the infinite-dimensional setting. This reformulation ensures consistent algorithmic behavior across various discretizations, as demonstrated in [22, 24, 15, 23, 44].
@@
To address these challenges, the Bayesianize-then-discretize approach has attracted significant attention from researchers in recent years [9, 14, 17].
@@
A primary challenge in Bayesian inference is the efficient extraction of posterior information. From the Bayesianize-then-discretize perspective, an infinite-dimensional Markov chain Monte Carlo (MCMC) algorithm known as preconditioned Crank-Nicolson (pCN) has been proposed and analyzed in detail [15, 37]. This algorithm maintains consistent sampling efficiency across different discretizations. In addition to the pCN algorithm, other types of infinite-dimensional sampling algorithms have been proposed, such as the infinite-dimensional sequential Monte Carlo algorithm [4] and the infinite-dimensional importance sampling algorithm [1]. To enhance sampling efficiency, infinite-dimensional MCMC algorithms with gradient and geometric informative proposals have been developed. Examples include the infinite-dimensional Metropolis-adjusted Langevin algorithm [10] and the geometric pCN algorithm [3]. Although these algorithms possess mesh independence properties, they are challenging to apply to large-scale inverse problems of PDEs, such as full waveform inversion [34].
@@
In the finite-dimensional setting, variational inference (VI) methods have been extensively studied in machine learning to mitigate the computational demands of Markov chain Monte Carlo (MCMC) sampling algorithms [49]. For example, a mean-field assumption based VI approach was employed to solve finite-dimensional inverse problems with hyper-parameters in prior and noise distributions [20, 25, 26]. Projected Stein variational gradient descent methods were constructed to solve inverse problems with low-intrinsic dimensions [13], and normalizing flow in Euclidean space have been extensively studied as a means to approximate target posterior distributions by transforming simple distributions [40, 46, 35].
@@
Here, we briefly outline the key distinctions between MCMC sampling and variational inference (see [6] for details). While MCMC methods typically require higher computational resources, they guarantee asymptotically exact samples from the target distribution [42]. In contrast, variational inference seeks an approximate distribution that is computationally tractable, leading to significant computational gains despite some loss of accuracy compared to MCMC. Variational inference’s compatibility with optimization techniques such as stochastic gradient descent and distributed optimization [41, 31, 48] makes it well-suited for large datasets and rapid model exploration. MCMC excels in scenarios with smaller datasets where precise sampling is prioritized. Another critical factor influencing the choice between MCMC and variational inference is the geometry of the posterior distribution. For instance, mixture models often exhibit multiple modes corresponding to different label permutations. In such cases, Gibbs sampling excels due to its ability to efficiently explore these modes [5]. However, when Gibbs sampling is not feasible, variational inference can outperform general MCMC methods, such as Hamiltonian Monte Carlo, even for smaller datasets [30].
@@
In contrast to the extensive research on finite-dimensional problems, the application of VI methods to infinite-dimensional settings remains relatively unexplored. Generally speaking, there are two main research directions for VI methods in infinite-dimensional settings:
@@
    Parameterize measures directly: This involves directly parameterizing the approximate measure and optimizing these parameters to approximate the posterior. For instance, when restricting approximations to Gaussian measures, a novel Robbins-Monro algorithm was developed using a calculus-of-variations perspective [38, 39]. Under the classical mean-field assumption, a general VI framework within separable Hilbert spaces was recently proposed [24, 23, 44].
@@
    Parameterize transformation: By parameterizing a transformation, a simple measure can be mapped into a more complex measure to approximate the posterior. Optimization of these transformation parameters drives the approximation process. For example, the infinite-dimensional Stein variational gradient descent introduced in [21] represents a function space particle optimization method with rigorous mathematical underpinnings in separable Hilbert spaces.
@@
In this work, we focus on defining normalizing flow in infinite-dimensional spaces, aligning with the second research direction. Specifically, we derive general transformation forms capable of mapping simple measures to complex distributions within the function space. When constructing the normalizing flows based infinite-dimensional variational inference (NF-iVI) algorithm, it is crucial to ensure that the post-transformed measure is absolutely continuous with respect to the prior measure. Otherwise, the Radon-Nikodym (RN) derivative of the post-transformed measure with respect to the prior measure will become meaningless, which would contradict the entire theoretical framework of the infinite-dimensional variational inference algorithm [39, 23]. Leveraging the properties of Gaussian measures [7], we develop a rigorous theoretical framework for transformations within functional normalizing flow. By formulating specific conditions, our approach guarantees the equivalence between the post-transformed and pre-transformed measures (detailed in Subsection 2.2), ensuring the model’s well-definedness in function space.
@@
In addition, we have devised a strategy to alleviate the computational demands of NF-iVI. Variational inference algorithms require analyzing the Kullback-Leibler (KL) divergence between the approximate and target posterior measures, which involves taking an expectation over the approximate measure. This typically necessitates computationally intensive Monte Carlo estimation. In our algorithm, the derivative of the KL divergence with respect to the parameters of the approximate measure must be calculated at each iteration. This necessitates computing the KL divergence and its gradient at each iteration, significantly increasing computational cost. While the NF-iVI algorithm offers advantages over traditional methods like pCN, it still presents considerable computational challenges (see Section 4 for details). We observe that for a fixed inverse problem model, distinct measurement data induce distinct posterior distributions, necessitating computationally intensive model retraining for each new set of measurement data. To enhance the computational efficiency of our algorithm when processing new measurement data, we propose to incorporate a conditional component into functional normalizing flow. By introducing measurement data as a conditional input, we enable the model to adapt to different measurement data without requiring extensive retraining. When the conditional network is trained, we can directly obtain an acceptable estimate of the posterior for any measurement data. We refer to this approach conditional normalizing flows based infinite-dimensional variational inference (CNF-iVI). Furthermore, if we want to obtain a more accurate approximation, we can consider further training for specific measurement data based on the pre-trained conditional functional normalizing flow.
@@
    We have applied NF-iVI and CNF-iVI to typical linear and nonlinear inverse problems, specifically elliptic inverse problems and steady-state Darcy flow inverse permeability estimations, thereby validating their efficiency.
The outline of this paper is as follows: In Section 2, we briefly introduce the theoretical framework and algorithmic process of NF-iVI. In Subsection 2.1, we present the infinite-dimensional Bayesian theory proposed in [38, 39] and provide a brief overview of infinite-dimensional variational inference algorithms. In Subsection 2.2, we introduce the NF-iVI algorithm in function spaces. In Subsections 2.3 and 2.4, we provide four examples of functional normalizing flow. In Subsection 2.5, we prove the discretization-invariance of the proposed functional normalizing flows. In Section 3, based on NF-iVI, we introduce the algorithmic process of CNF-iVI. In Subsection 3.1, we construct an additional conditional input component, allowing the parameters of functional normalizing flow to be determined by observational information. We also present the detailed training procedure for CNF-iVI. In Subsection 3.2, we describe a method to improve the results of conditional functional normalizing flow. In Sections 4 and 5, we apply the algorithms to two typical inverse problems, verifying the feasibility of both the NF-iVI and CNF-iVI algorithms. Additionally, in each numerical simulation, we demonstrate the mesh independence as expected for the Bayesianize-then-discretize approach. In Section 6, we summarize our achievements, acknowledge some deficiencies, and explore further research directions.
@@ 2 Functional Normalizing Flow
In this section, we establish the theoretical framework of functional normalizing flow. To circumvent the obstacle posed by the singularity of measures in infinite-dimensional spaces, we introduce a novel theoretical framework. Building upon this foundation, four specific flow models are proposed and shown to satisfy the established theoretical criteria.
@@ 2.1 Bayesian Approach and Variational Inference
In this subsection, we introduce the foundational concepts of infinite-dimensional Bayesian inverse problems and the underlying principles of infinite-dimensional variational inference.
Let $\mathcal{H}_{u}$, $\mathcal{H}_{w}$ be separable Hilbert spaces representing the parameter space and solution space, respectively, and $N_{d}$ be a positive integer. Denote that $\mathcal{N}(u,\mathcal{C})$ is a Gaussian measure with mean $u$ and covariance operator $\mathcal{C}$. The inverse problem can be described as follow:
@@
where $\bm{d}\in\mathbb{R}^{N_{d}}$ is the measurement data, $u\in\mathcal{H}_{u}$ is the interested parameter, $\mathcal{G}$ is the PDE solving operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{w}$, $\mathcal{S}$ is the observation operator from $\mathcal{H}_{w}$ to $\mathbb{R}^{N_{d}}$, and $\bm{\epsilon}$ is a Gaussian random vector with zero mean and covariance matrix $\bm{\Gamma}_{\text{noise}}:=\tau^{-1}\textbf{I}$ ($\tau$ is a fixed positive number, and I denotes the $N_{\bm{d}}$-dimensional identity matrix), which means
@@
Based on the framework of infinite-dimensional Bayesian inference [47], we are able to preserve the fundamental Bayes’ formula for the inverse problem:
@@
where $\Phi:\mathcal{H}_{u}\rightarrow\mathbb{R}$ is defined as
$\Phi(u)=\frac{1}{2}\lVert\bm{d}-\mathcal{S}\mathcal{G}(u)\rVert^{2}_{\bm{\Gamma}_{\text{noise}}},$ (2.4)
where $\lVert\cdot\rVert_{\bm{\Gamma}_{\text{noise}}}:=\lVert\bm{\Gamma}_{\text{noise}}^{-1/2}\cdot\rVert$ and $Z_{\mu}$ is a positive finite constant given by
@@
Directly extracting quantitative information from the posterior distribution in Bayes’ formula (2.3) is computationally prohibitive [15]. To circumvent this challenge, variational inference approximates the intractable posterior with a more tractable measure. We select a set of measures $\mathcal{M}(\mathcal{H}_{u})$ on space $\mathcal{H}_{u}$ as the approximating measure set. To ensure that the Radon-Nikodym (RN) derivative between measures is meaningful, we require that any measure within $\mathcal{M}(\mathcal{H}_{u})$ is equivalent to the prior measure $\mu_{0}$ [39]. For any $\nu\in\mathcal{M}(\mathcal{H}_{u})$, the KL divergence between $\nu$ and posterior $\mu$ is given by
@@
When solving the minimization problem (2.5), it will usually be convenient to parameterize the set $\mathcal{M}(\mathcal{H}_{u})$ in a suitable way: Overly intricate measures can hinder optimization, while overly simplistic ones may compromise approximate accuracy. To address this trade-off, [39, 38] employed carefully designed Gaussian measures, whereas [24, 44, 23] opted for the mean-field approach.

### v2

@@ Abstract
Inverse problems of partial differential equations are ubiquitous across various scientific disciplines and can be formulated as statistical inference problems using Bayes’ theorem. To address large-scale problems, it is crucial to develop discretization-invariant algorithms, which can be achieved by formulating methods directly in infinite-dimensional space. We propose a novel normalizing flow based infinite-dimensional variational inference method (NF-iVI) to extract posterior information efficiently. Specifically, by introducing well-defined transformations, the prior in Bayes’ formula is transformed into post-transformed measures that approximate the posterior. To circumvent the issue of mutually singular probability measures, we formulate general conditions for the employed transformations. As guiding principles, these conditions yield four concrete transformations. Additionally, to minimize computational demands, we have developed a conditional normalizing flow variant, termed CNF-iVI, which is adapt at processing measurement data of varying dimensions while requiring minimal computational resources. We apply the proposed algorithms to three typical inverse problems governed by the simple smooth equation, the steady-state Darcy flow equation, and the electric impedance tomography. Numerical results confirm our theoretical findings, illustrate the efficiency of our algorithms, and verify the discretization-invariant property.
@@ 1 Introduction
Driven by their widespread applications in seismic exploration, radar imaging, and other fields, inverse problems involving partial differential equations (PDEs) have witnessed significant advancements in recent decades Kirsch2011Book, Benning2018ActaNum. As computational power continues to grow, researchers are increasingly focused on not only obtaining estimated solutions but also conducting statistical analyses to quantify uncertainties, which is crucial for tasks such as artifact detection Zhou2020SJIS. The Bayesian inverse approach offers a robust framework for addressing inverse problems within the context of PDEs by transforming them into statistical inference problems, thereby enabling the analysis of parameter uncertainties dashti2013bayesian.
Typically, inverse problems of PDEs are posed in infinite-dimensional spaces Kirsch2011Book, which presents challenges for applying well-established finite-dimensional Bayesian inference techniques bishop2006pattern, kaipio2006statistical. To bridge this gap, two predominant strategies have been developed:
@@
    Discretize-then-Bayesianize: PDEs are initially discretized to approximate the original problem in some finite-dimensional space, and the reduced, approximated problem is then solved using Bayes’ method kaipio2006statistical.
@@
    Bayesianize-then-discretize: Bayes’ formula and algorithms are initially constructed in infinite-dimensional space, and after the infinite-dimensional algorithm is built, some finite-dimensional approximation is carried out dashti2013bayesian.
Both approaches offer distinct advantages and disadvantages. The Discretize-then-Bayesianize approach allows us to employ all the Bayesian inference methods developed in the statistical literature kaipio2006statistical, Calvetti2023Book to solve the inverse problems. However, due to the infinite-dimensional nature of the original problems, two critical challenges arise:
@@
    Model consistency: Finite-dimensional models, such as those representing prior probability measures, exhibit distinct properties compared to their infinite-dimensional counterparts. A notable example is the total variation prior, which has been extensively studied in lassas2004can, Dashti2012IPI.
@@
    Algorithm applicability: To preserve the intrinsic structure of original infinite-dimensional problems, algorithms initially designed for finite-dimensional spaces must be carefully adapted to the infinite-dimensional setting. This reformulation ensures consistent algorithmic behavior across various discretizations, as demonstrated in Jia2021IPI, jia2021variational, cotter2013, Jia2022VINet, sui2024noncentered.
@@
To address these challenges, the Bayesianize-then-discretize approach has attracted significant attention from researchers in recent years bui2013computational, Cotter2009IP, dashti2013bayesian.
@@
A primary challenge in Bayesian inference is the efficient extraction of posterior information. From the Bayesianize-then-discretize perspective, an infinite-dimensional Markov chain Monte Carlo (MCMC) algorithm known as preconditioned Crank-Nicolson (pCN) has been proposed and analyzed in detail cotter2013, Pillai2014SPDE. This algorithm maintains consistent sampling efficiency across different discretizations. In addition to the pCN algorithm, other types of infinite-dimensional sampling algorithms have been proposed, such as the infinite-dimensional sequential Monte Carlo algorithm Beskos2015SC and the infinite-dimensional importance sampling algorithm Agapiou2017SS. To enhance sampling efficiency, infinite-dimensional MCMC algorithms with gradient and geometric informative proposals have been developed. Examples include the infinite-dimensional Metropolis-adjusted Langevin algorithm Thanh2016IPI and the geometric pCN algorithm Beskos2017JCP. Although these algorithms possess mesh independence properties, they are challenging to apply to some large-scale inverse problems of PDEs Fichtner2011Book.
@@
In the finite-dimensional setting, variational inference (VI) methods have been extensively studied in machine learning to mitigate the computational demands of MCMC sampling algorithms zhang2018advances. For example, a mean-field assumption based VI approach was employed to solve finite-dimensional inverse problems with hyper-parameters in prior and noise distributions Guha2015JCP, Jin2012JCP, jin2010hierarchical. Projected Stein variational gradient descent methods were constructed to solve inverse problems with low-intrinsic dimensions Chen2021SISC, and normalizing flow in Euclidean space have been extensively studied as a means to approximate target posterior distributions by transforming simple distributions planarflow, sylvesterflow, householderflow1, MR4749392.
@@
Here, we briefly outline the key distinctions between MCMC sampling and variational inference (see blei2017variational for details). While MCMC methods typically require higher computational resources, they guarantee asymptotically exact samples from the target distribution robert2004monte. In contrast, variational inference seeks an approximate distribution that is computationally tractable, leading to significant computational gains despite some loss of accuracy compared to MCMC. Variational inference’s compatibility with optimization techniques such as stochastic gradient descent and distributed optimization robbins1951stochastic, kushner2003stochastic, welling2011bayesian makes it well-suited for large datasets and rapid model exploration. MCMC excels in scenarios with smaller datasets where precise sampling is prioritized. Another critical factor influencing the choice between MCMC and variational inference is the geometry of the posterior distribution. For instance, mixture models often exhibit multiple modes corresponding to different label permutations. In such cases, Gibbs sampling excels due to its ability to efficiently explore these modes bishop2006pattern. However, when Gibbs sampling is not feasible, variational inference can outperform general MCMC methods, such as Hamiltonian Monte Carlo, even for smaller datasets kucukelbir2017automatic.
@@
In contrast to the extensive body of research in finite-dimensional settings, the application of variational inference methods to infinite-dimensional settings remains relatively unexplored. Generally speaking, there are two main research directions for VI methods in infinite-dimensional settings:
@@
    Parameterize measures directly: This involves directly parameterizing the approximate measure and optimizing these parameters to approximate the posterior. For instance, when restricting approximations to Gaussian measures, a novel Robbins-Monro algorithm was developed using a calculus-of-variations perspective pinski2015algorithms, Pinski2015SIAMMA. Under the classical mean-field assumption, a general VI framework within separable Hilbert spaces was recently proposed jia2021variational, Jia2022VINet, sui2024noncentered.
@@
    Parameterize transformation: By parameterizing a transformation, a simple measure can be mapped into a more complex measure to approximate the posterior. Optimization of these transformation parameters drives the approximation process. For example, the infinite-dimensional Stein variational gradient descent introduced in jia2021stein represents a function space particle optimization method with rigorous mathematical underpinnings in separable Hilbert spaces.
@@
In this work, we focus on defining normalizing flow in infinite-dimensional spaces, aligning with the second research direction. Specifically, we derive general transformation forms capable of mapping simple measures to complex distributions within the function space. It is noteworthy that the definition of normalizing flow in function space has been considered before in shiuniversal. However, the objectives of this work differ significantly from ours. In shiuniversal, the authors focus on constructing a generative model to fit a collection of functions $\mathbb{D}=\{u_{i}\}_{i=1}^{N}$ resides in some function spaces, rather than approximating the posterior of Bayes’ formula. More importantly, our model establishes a rigorous theory guaranteeing the equivalence of the measure before and after the transformation, which provides a solid theoretical foundation for posterior approximation. Precisely, it is crucial to ensure that the post-transformed measure is absolutely continuous with respect to the prior measure when constructing the normalizing flow-based infinite-dimensional variational inference (NF-iVI) algorithm. Failure to ensure this condition would render the Radon-Nikodym (RN) derivative of the post-transformed measure with respect to the prior measure meaningless, thus contradicting the entire theoretical framework of the infinite-dimensional variational inference algorithm Pinski2015SIAMMA, Jia2022VINet. Leveraging the properties of Gaussian measures gaussianmeasure, we develop a rigorous theoretical framework for transformations within functional normalizing flow. By formulating specific conditions, our approach guarantees the equivalence between the post-transformed and pre-transformed measures (detailed in Subsection 2.2), ensuring the model’s well-definedness in function space.
@@
In addition, we have devised a strategy to alleviate the computational demands of NF-iVI. Variational inference algorithms require analyzing the Kullback-Leibler (KL) divergence between the approximate and target posterior measures, which involves taking an expectation over the approximate measure. This typically necessitates computationally intensive Monte Carlo estimation. In our algorithm, the derivative of the KL divergence with respect to the parameters of the approximate measure must be calculated at each iteration. This necessitates computing the KL divergence and its gradient at each iteration, significantly increasing computational cost. While the NF-iVI algorithm offers advantages over traditional methods like pCN, it still presents considerable computational challenges (see Section 4 for details). We observe that for a fixed inverse problem model, distinct measurement data induce distinct posterior distributions, necessitating computationally intensive model retraining for each new set of measurement data. To enhance the computational efficiency of our algorithm when processing new measurement data, we propose to incorporate a conditional component into functional normalizing flow. By introducing measurement data as a conditional input, we enable the model to adapt to different measurement data without requiring extensive retraining. When the conditional network is trained, we can directly obtain an acceptable estimate of the posterior for any measurement data. We refer to this approach conditional normalizing flow based infinite-dimensional variational inference (CNF-iVI). Furthermore, if we want to obtain a more accurate approximation, we can consider further training for specific measurement data based on the pre-trained conditional functional normalizing flow.
@@
    We have applied NF-iVI and CNF-iVI to typical linear and nonlinear inverse problems, specifically elliptic inverse problems, steady-state Darcy flow inverse permeability estimations and electric impedance tomography, thereby validating their efficiency.
The outline of this paper is as follows: In Section 2, we briefly introduce the theoretical framework and algorithmic process of NF-iVI. In Subsection 2.1, we present the infinite-dimensional Bayesian theory proposed in pinski2015algorithms, Pinski2015SIAMMA and provide a brief overview of infinite-dimensional variational inference algorithms. In Subsection 2.2, we introduce the NF-iVI algorithm in function spaces. In Subsection 2.3, we provide a detailed comparison between the conventional neural operator and the functional normalizing flow. In Subsections 2.4 and 2.5, we provide four examples of functional normalizing flow. In Subsection 2.6, we prove the discretization-invariance of the proposed functional normalizing flows. In Section 3, based on NF-iVI, we introduce CNF-iVI, adding an auxiliary conditional input to enable the functional normalizing flow parameters to be determined by the measurement information. The detailed training procedure for CNF-iVI is also provided. Furthermore, we describe a method to improve the results of conditional functional normalizing flow. In Sections 4 and 5, we apply the algorithms to three typical inverse problems, verifying the feasibility of both the NF-iVI and CNF-iVI algorithms. Additionally, in each numerical simulation, we demonstrate the mesh independence as expected for the Bayesianize-then-discretize approach. In Section 6, we summarize our achievements, acknowledge some deficiencies, and explore further research directions.
@@ 2 Functional Normalizing Flow
In this section, we establish the theoretical framework of the functional normalizing flow and demonstrate its specific application to the PDE inverse problem involving infinite-dimensional parameters and finite-dimensional measurements. To circumvent the obstacle posed by the singularity of measures in infinite-dimensional spaces, we introduce a novel theoretical framework. Building upon this foundation, four specific flow models are proposed and shown to satisfy the established theoretical criteria.
@@ 2.1 Bayesian Approach and Variational Inference
For a large class of inverse problems involving partial differential equations, sparse measurement data are practically adopted Cotter2009IP, nickl2020bernstein, as such data are more easily acquired. In this subsection, we introduce the foundational concepts of infinite-dimensional Bayesian inverse problems with finite-dimensional sparse data, as well as the underlying principles of infinite-dimensional variational inference.
Let $\mathcal{H}_{u}$, $\mathcal{H}_{w}$ be separable Hilbert spaces representing the parameter space and solution space, respectively, and $N_{d}$ be a positive integer. Denote $\mathcal{N}(u,\mathcal{C}_{0})$ as a Gaussian measure with mean $u$ and covariance operator $\mathcal{C}_{0}$. The inverse problem can be described as
@@
where $\bm{d}\in\mathbb{R}^{N_{d}}$ is the measurement data, $u\in\mathcal{H}_{u}$ is the parameter of interest, $\mathcal{G}$ is the PDE solution operator from $\mathcal{H}_{u}$ to $\mathcal{H}_{w}$, $\mathcal{S}$ is the measurement operator from $\mathcal{H}_{w}$ to $\mathbb{R}^{N_{d}}$, and $\bm{\epsilon}$ is a Gaussian random vector with zero mean and covariance matrix $\bm{\Gamma}_{\text{noise}}\mathrel{\mathop{\ordinarycolon}}=\tau^{-1}\textbf{I}$ ($\tau$ is a fixed positive number, and I denotes the matrix of identity), which means
@@
Based on the framework of infinite-dimensional Bayesian inference weglein2003inverse, we are able to preserve the fundamental Bayes’ formula for the inverse problem:
@@
where $\Phi\mathrel{\mathop{\ordinarycolon}}\mathcal{H}_{u}\rightarrow\mathbb{R}$ is defined as
$\Phi(u)=\frac{1}{2}\lVert\bm{d}-\mathcal{S}\mathcal{G}(u)\rVert^{2}_{\bm{\Gamma}_{\text{noise}}}$ (2.4)
with $\lVert\cdot\rVert_{\bm{\Gamma}_{\text{noise}}}\mathrel{\mathop{\ordinarycolon}}=\lVert\bm{\Gamma}_{\text{noise}}^{-1/2}\cdot\rVert$, and $Z_{\mu}$ is a positive finite constant given by
@@
Directly extracting quantitative information from the posterior distribution in Bayes’ formula (2.3) is computationally prohibitive cotter2013. To circumvent this challenge, variational inference approximates the intractable posterior with a more tractable measure. We select $\mathcal{M}(\mathcal{H}_{u})$, a set of measures on the space $\mathcal{H}_{u}$, as the approximating measure set. To ensure that the Radon-Nikodym (RN) derivative is well-defined, every measure in $\mathcal{M}(\mathcal{H}_{u})$ is required to be equivalent to the prior measure $\mu_{0}$ Pinski2015SIAMMA.
For any $\nu\in\mathcal{M}(\mathcal{H}_{u})$, the KL divergence between $\nu$ and posterior $\mu$ is given by
@@
Solving the minimization problem (2.5) typically requires a suitable parameterization of the set $\mathcal{M}(\mathcal{H}_{u})$: Overly intricate measures can hinder optimization, while overly simplistic ones may compromise approximate accuracy. To address this trade-off, Pinski2015SIAMMA, pinski2015algorithms employed carefully designed Gaussian measures, whereas jia2021variational, sui2024noncentered, Jia2022VINet opted for the mean-field approach.